For the following situations, calculate the cash surplus or deficit. Cash Inflows Cash Outflows Difference Surplus/Deficit 3,540S 4,788 $ 4,387 S 3,238 4,855 4,198
Answer #1

If cash inflows is greater than cash outflows;there is a surplus or else there is a deficit.

cash inflows cash outflows Difference(cash inflows-cash outflows) Surplus/Deficit
3540 3238 (3540-3238)=$302 Surplus
4788 4855 (4788-4855)=$67(Negative). Deficit
4387 4198 (4387-4198)=$189 Surplus